OpenStudy (anonymous):

Please help! Factor Completely: x^2-2x-8 I have the two numbers that have the product of -8 and sum of -2. -4 and 2 But I can't get past that

OpenStudy (anonymous):

the way that I used to think of these when I first learned them was to set up the brackets and input them until I got the right answer. you have done all of the hard work. put this down with blanks then fill in until the answer works. (x )(x ) Since you already have the -4, 2 just put those in, and check by multiplying out (x-4 )(x+2 ) = x^2 +2x-4x-8 = x^2-2x-8

OpenStudy (anonymous):

Awesome! Thank you :D